A wire grating is made of 200 wires per cm placed parallel to each other with uniform separation . The diameter of each wire is 0.025mm calculate the angle of diffraction for third order spectrum and also find the absent spectra if any, wavelenght of light used is 6000AO.

Grade:12th Pass

1 Answers

suresh kumar
3 years ago
each wire diameter is 0.025mm,
there are 200 wires in 1cm
so, the total width of 200 wires = 200*0.025mm = 5mm per 1cm (since there are 200 wires in 1cm)
remaining space is the separation.
hence the separation between two wires = 5mm/200 = 0.025 mm
we know dsinθ = mλ
   here m=3 , d = 0.025mm and θ = θ3 we need to find θ3
     0.025mm sinθ3 = 3*6000A
                on sloving we get θ3 = sin-1 (0.072)